Hi, I have no idea on how to do solve this word problem:
The Rent-Me Car Rental charges $15 per day plus $.22 per mile to rent a car. Determine a linear function that can be used to calculate daily car rentals. Then use that function to determine the cost of renting a car for a day and driving 300 miles.
Please help me:) Thanks

Found 2 solutions by checkley71, mathslover:
Answer by checkley71(8403)

FORMULA IS:15*D+.22*M WHERE D=NUMBER OF DAYS & M=NUMBER OF MILES.
THUS THE COST FOR 1 DAY RENTAL WITH 300 MILES IS:
15+.22*300=15+66=$81 ANSWER FOR 1 DAYS RENTAL WITH 300 MILES.

Answer by mathslover(157)

Let "m" be the number of miles driven in day
Given Charges per mile = $.22
Charges for driving m miles = $(.22*m)
Total rental for the day is $(15 + .22m) where m is thenumber of miles driven in a day
Cost when 300 miles are driven in a day = $(15 + .22*300)
=$(15+ 66)
=$81